OpenStudy (anonymous):

The temperature of 6.24 L of a gas is increased from 25.0°C to 55.0°C at constant pressure. The new volume of the gas is A. 13.7 L. B. 5.67 L. C. 6.87 L. D. 2.84 L.

OpenStudy (anonymous):

@thomaster please help?

OpenStudy (anonymous):

Let me try this was our lesson last month....

OpenStudy (anonymous):

@EngrChong i think its (C)

OpenStudy (anonymous):

Do you have a scientific calculator....?

OpenStudy (anonymous):

i do not

OpenStudy (anonymous):

I have....I'll solve it....

OpenStudy (anonymous):

Charles' law

OpenStudy (anonymous):

V/T=K

OpenStudy (jfraser):

to use the gas laws, you have to use the temperature in Kelvin, not Celsius or Farenheit

OpenStudy (anonymous):

You teach her @JFraser .....Im not so good at explaining....

OpenStudy (anonymous):

K= C* + 273

OpenStudy (anonymous):

how would this situation play out? a demonstration? a drawing?

OpenStudy (anonymous):

K= 25.0 C* +273 K= 55.0 C* +273

OpenStudy (anonymous):

Can you do it from here?

OpenStudy (anonymous):

V1/T1=V2/T2

OpenStudy (anonymous):

would you divide them ? its been a year since i took this class

OpenStudy (jfraser):

You have 3 of the 4 pieces you need: V1, T1, and T2. Rearrange the equation and solve for T2, then plug in the pieces and solve

OpenStudy (abmon98):

At Constant Pressure- When we increase the temperature, the molecules move more rapidly, gas expands as it occupies more volume. Use Charles Law to find V2. V1/T1=V2/T2 Note: Temperature is measured in kelvin so add the values of temperature to 273

OpenStudy (anonymous):

(V1)T2/T1=V2

OpenStudy (anonymous):